# Control-Aware Trajectory Predictions for Communication-Efficient Drone Swarm Coordination in Cluttered Environments

### v1

@@ III Control-aware learning-based trajectory prediction
To address the challenge imposed by limited communication, we devise a learning-based approach to predict the trajectories of surrounding UAVs of each UAV $i\in\mathcal{V}$. Here, we seek to predict the trajectories of surrounding UAVs by estimating their planned trajectories at the previous time step. The reason is three-fold. First, due to embedded motion control modules in UAVs, we envision the planned trajectory to be close to the actual trajectory, especially if the planning horizon is relatively short. Second, the estimation of the planned trajectories enables us to encode more physical information (e.g., the decision-making process of UAVs) and thus can be more sample efficient in training. Third, using the planned trajectories to represent future trajectories is commonly seen in existing literature [44, 28, 8, 33, 43], which enables us to better integrate our approach with state-of-the-art control algorithms. Mathematically, let $\tilde{\bm{U}}_{ij}^{k-1}$ be the planned trajectory of UAV $j$ calculated at time step $k-1$ and estimated by UAV $i$ at the beginning of time step $k$. For brevity, We simplify the notion $\tilde{\bm{U}}_{ij}^{k-1}$ to $\tilde{\bm{U}}_{j}^{k-1}$ by omitting $i$ if there is no ambiguity. Hence, the predicted trajectory of all neighbors of UAV $i$, $\tilde{\bm{U}}_{-i}^{k}$, can be represented by $\tilde{\bm{U}}_{-i}^{k}=\{\tilde{\bm{U}}_{j}^{k-1}\}_{j\in\mathcal{N}_{i}}$.
@@
The proposed trajectory prediction approach involves four components as illustrated in Figure 2: (i) a Variational Auto-Encoder (VAE)-based component that allows each UAV to compress its planned trajectory and reconstruct the compressed messages sent by other UAVs, (ii) an EvolveGCN (EG)-based trajectory prediction component for the estimation of the planned trajectory in scenarios with dynamic sensing/communication graphs, (iii) a Bayesian layer to integrate the reconstructed planned trajectories by VAE and the estimated planned trajectories by the trajectory prediction component, and (iv) a KKT-enhanced training approach that devises a KKT-based differential layer to encode the structural information of the embedded QP into the trained neural network. The details of these components are detailed below in Sections III-AIII-D. Note that the developed learning-based model is implemented in each UAV with identical parameters.
III-A VAE for trajectory compression and reconstruction
A VAE is designed for trajectory compression and reconstruction, which consists of an encoder $\phi_{\rm{e}}$ and a decoder $\phi_{\rm{d}}$. The encoder is leveraged by UAV $j\in\mathcal{N}_{i}$ to compress its planned trajectory $\bm{U}_{j}^{k-1}\in\mathbb{R}^{3P}$ into a low-dimensional message $\bm{z}_{j}^{k}\in\mathbb{R}^{L}$ ($L<3P$). Specifically, the encoder is implemented as a neural network that outputs a vector of means $\phi_{\rm{e}}(\bm{U}_{j}^{k-1})\in\mathbb{R}^{L}$ and standard deviations $\bm{\eta}_{j}^{k}\in\mathbb{R}^{L}$ of independent Gaussian distributions, from which the message $\bm{z}_{j}^{k}$ is sampled and transmitted to UAV $i$. After UAV $i$ receives the message $\bm{z}_{j}^{k}$, it uses the decoder to reconstruct an estimated planned trajectory $\tilde{\bm{U}}_{j,\text{VAE}}^{k-1}=\phi_{\rm{d}}(\bm{z}_{j}^{k})$. In this paper, both the encoder network $\phi_{\rm{e}}$ and decoder network $\phi_{\rm{d}}$ are implemented as fully-connected layers. In addition, the trajectory reconstructed by the decoder naturally makes VAE an estimator of the true trajectory.
III-B EvolveGCN (EG)-based trajectory prediction
@@
Our EvolveGCN (EG)-based trajectory prediction component enables UAV $i$ to estimate the planned trajectories of all its neighbors, $\tilde{\bm{U}}_{j,\rm{EG}}^{k-1}$, leveraging (i) the geometric information of obstacles and (ii) the location information within a horizon of $H$ past time steps $\bm{p}^{k,\text{hist}}=\{\bm{p}_{j}^{k-h}\}_{j\in\mathcal{N}_{i},h\in\{0,1,\cdots,H\}}$ of surrounding UAVs, measured by the onboard sensors of UAV $i$, and (iii) the planned trajectories of UAV $\tilde{\bm{U}}_{j}^{k-2}$ estimated at the previous time step $k-1$.
@@
As shown in Figure 2, the overall architecture involves four parts: (i) a Long Short-Term Memory (LSTM)-based encoder to characterize the correlations between planned trajectories estimated at consecutive time steps, (ii) an obstacle encoder to extract the geometric features of obstacles, (iii) an EvolveGCN layer [7] to extract the spatiotemporal correlations of UAVs on the dynamic sensing graph, and (iv) an LSTM-based decoder to generate estimated $\tilde{\bm{U}}_{j,EG}^{k-1}$ and its standard deviation $\bm{\sigma}_{j,\text{EG}}^{k}$. Parts (i), (ii), and (iv) build on the trajectory planning neural network developed in [8] and extend it by allowing a probabilistic output representation for the integration with VAE reconstruction outcomes. Moreover, unlike [8] that requires each UAV to measure the real-time locations of all other UAVs, we consider a dynamic sensing graph evolving with the movement of the swarm to capture the topological information of the UAV swarm. To this end, we add Part (iii) with an EvolveGCN (EG) architecture, which represents the swarm alignment by a graph convolutional network (GCN) and evolves its parameters along the temporal dimension using LSTM without resorting to node embeddings.
The details of the four parts are described as follows.
@@
(i) LSTM-based encoder. An LSTM layer is leveraged to exploit the temporal correlations between the planned trajectories of UAV $j$ at consecutive time steps. This encoder can be mathematically represented as:
@@
$\bm{y}_{j}^{k}=\text{LSTM}_{\text{query}}\left(\tilde{\bm{U}}_{j}^{k-2}\right),\leavevmode\nobreak\ j\in\mathcal{N}_{i}$ (9)
@@
(ii) Obstacle encoder. Since we only consider static obstacles, we design the obstacle encoder as a fully connected layer with an input of the parameters characterizing the obstacles. This encoder can be mathematically represented as:
$\bm{o}^{k}=\text{FC}_{\text{obs}}\left(\bm{C}_{m}\right),\leavevmode\nobreak\ {m\in M}$ (10)
@@
(iii) EG layers. As mentioned above, the EG layers are adopted to extract the spatiotemporal correlations of UAVs in time-varying sensing graphs. The input to the EG layers includes the position information of $H$ past time steps $\bm{p}^{k,\rm{hist}}$ within the sensing range of UAV $i$. The output of the EG layers, $\bm{g}^{k}$, serves as an input to the LSTM decoder, i.e., Part (iv).
Each EG layer involves a GCN layer and an LSTM layer, whereby the GCN layer captures the topology of the sensing graph seen by UAV $i$, represented as adjacency matrix $\bm{G}_{i}^{k}\in\mathbb{R}^{n\times n}$, where the elements involving unseen UAVs are set to zero. The LSTM layer is employed to model the evolution of the GCN weight matrix as a dynamical system. Specifically, let $\bm{W}^{k,(l)}$ denote the dynamic GCN weight matrix of the $l$-th layer at time step $k$, and $\bm{H}^{k,(l)}$ represents the node embeddings of the $l$-th layer at time step $k$. The EG layer can be written as:
$\bm{H}^{k+1,(l+1)}$ $=\text{GCONV}(\bm{G}_{i}^{k},\bm{H}^{k,(l)},\bm{W}^{k,(l)})$ (11)
$\bm{W}^{k+1,(l)}$ $=\text{LSTM}_{\text{eg}}\left(\bm{W}^{k,(l)}\right).$ (12)
where the initial embedding matrix $\bm{H}^{k,(0)}$ includes the node features as the output of the LSTM-based encoder (i.e., $\bm{y}^{k}$), and the final embedding matrix $\bm{H}^{k,(L)}$ is the output of the EG layer (i.e., $\bm{g}^{k}$). The activation function $\pi(\cdot)$ is chosen as ReLU for all but output layers and an identity function for the output layer.
In (12), the evolution of the weight matrix $\bm{W}^{k,(l)}$ is solely based on the weight matrix at the previous time step, which is referred to as EGCU-O. An alternative way of formulating this LSTM layer is to incorporate node embeddings $\bm{H}^{k,(l)}$, i.e., $\bm{W}^{k+1,(l)}=\text{LSTM}\left(\bm{H}^{k+1,(l)},\bm{W}^{k,(l)}\right)$, referred to as EGCU-H [7]. In this paper, we use EGCU-O for its simplicity.
@@
(iv) LSTM-based decoder. The decoder synthesizes the outputs of the LSTM encoder, the obstacle encoder and the EG layer with an LSTM layer followed by a fully connected layer to generate the sequence of estimated planned trajectories of UAV $j$, $\tilde{\bm{U}}_{j}^{k-1}$, which can be written as:
$\tilde{\bm{U}}_{j,\text{EG}}^{k-1},\bm{\sigma}_{j,\text{EG}}^{k-1}=\text{Decoder}_{\text{LSTM}}\left(\bm{y}_{j}^{k},\bm{o}^{k},\bm{g}^{k}\right)$ (13)
@@
where the distribution of the planned trajectories is given as a Gaussian distribution with mean $\tilde{\bm{U}}_{j,\rm{EG}}^{k-1}$ and covariance matrix $\Sigma^{k-1}_{j,\text{EG}}=\left(\text{diag}(\bm{\sigma}_{j,\text{EG}}^{k-1})\right)^{2}$. We choose the covariance matrix as a diagonal matrix, a typical treatment in existing works to simplify the underlying neural network.
III-C Bayesian integration of predicted planned trajectories
@@
Note that both the VAE and EG-based trajectory prediction components provide an estimation of the planned trajectories. The aim of this component is to combine them via a simple Bayesian approach. Specifically, we treat the outcome of the trajectory prediction module as the prior, because such estimation will always be available despite communication failure. Specifically, the prior distribution is characterized by a Gaussian distribution written as:
$\bm{U}_{j}^{k-1}|\tilde{\bm{U}}_{j,\text{EG}}^{k-1}\sim N\left(\tilde{\bm{U}}_{j,\text{EG}}^{k-1},\Sigma^{k-1}_{j,\text{EG}}\right)$ (14)
@@
where $\bm{U}_{j}^{k-1}$ represents the true planned trajectories made by UAV $j$ at time step $k-1$.
@@
If UAV $i$ receives a message $\bm{z}_{j}$ from UAV $j$ and reconstructs it, the reconstructed outcome, $\tilde{\bm{U}}_{j,\text{VAE}}^{k-1}$, will serve as an observation, i.e.,
@@
$\tilde{\bm{U}}_{j,\text{VAE}}^{k-1}|\bm{U}_{j}^{k-1}\sim N\left(\bm{U}_{j}^{k-1},\Sigma^{k-1}_{\text{VAE}}\right),\leavevmode\nobreak\ j\in\mathcal{N}_{i}$ (15)
@@
where the covariance matrix $\Sigma^{k-1}_{\text{VAE}}$ can be obtained empirically from a well-trained VAE.
@@
Combining (14) and (15), the posterior distribution can be derived using the Bayesian formula as in Eq.(16). Here, notice that $\tilde{\bm{U}}_{j,\text{VAE}}^{k-1}$ and $\tilde{\bm{U}}_{j,\text{EG}}^{k-1}$ are independent.
$p(\bm{U}_{j}^{k-1}|\tilde{\bm{U}}_{j,\text{VAE}}^{k-1},\tilde{\bm{U}}_{j,\text{EG}}^{k-1})\propto p(\bm{U}_{j}^{k-1}|\tilde{\bm{U}}_{j,\text{EG}}^{k-1})p(\tilde{\bm{U}}_{j,\text{VAE}}^{k-1}|\bm{U}_{j}^{k-1}).$ (16)
@@
The final estimation can be calculated via maximum a posteriori (MAP) as $\tilde{\bm{U}}_{j}^{k-1}=\arg\max_{\bm{U}_{j}^{k-1}}p(\bm{U}_{j}^{k-1}|\tilde{\bm{U}}_{j,\text{VAE}}^{k-1},\tilde{\bm{U}}_{j,\text{EG}}^{k-1})$.
III-D KKT-informed training of the trajectory prediction neural network
Sections III-AIII-C jointly define a learning-based framework for trajectory prediction with a neural network model. This neural network is trained with data generated from simulations. For details about the generation of training data used in this paper, please refer to Section IV-B. One classical method to train the neural network is to minimize the $l_{2}$-loss, written as
$\mathcal{L}_{2}(\tilde{\bm{U}}_{j}^{k-1},\bm{U}_{j}^{k-1})=||\tilde{\bm{U}}_{j}^{k-1}-\bm{U}_{j}^{k-1}||_{2}^{2}$ (17)

### v2

@@ III Control-aware learning-based trajectory prediction
Figure 3: Example of KKT-informed training process, whereby UAV 1 wants to estimate the planned trajectories of UAV 2. UAV 1, 3, 4, and 5 are the closest neighbors of UAV 2.
@@
To address the challenge imposed by limited communication, we devise a learning-based approach to predict the trajectories of the surrounding UAVs of each UAV $i\in\mathcal{V}$. Here, we seek to predict the trajectories of surrounding UAVs by estimating their planned trajectories at the current time step. The reason is three-fold. First, due to embedded motion control modules in UAVs, we envision the planned trajectory to be close to the actual trajectory, especially if the planning horizon is relatively short. Second, the estimation of the planned trajectories enables us to encode more physical information (e.g., the decision-making process of UAVs) and thus can be more sample-efficient in training. Third, using the planned trajectories to represent future trajectories is commonly seen in existing literature [38, 17, 21, 30, 39], which enables us to better integrate our approach with state-of-the-art control algorithms. Mathematically, let $\tilde{\bm{U}}_{ij}^{k}$ be the predicted trajectory of UAV $j$ calculated by UAV $i$ at the beginning of time step $k$. For brevity, we simplify the notion $\tilde{\bm{U}}_{ij}^{k}$ to $\tilde{\bm{U}}_{j}^{k}$ by omitting $i$ if there is no ambiguity. Therefore, the predicted trajectory of all neighbors of the UAV $i$, $\tilde{\bm{U}}_{-i}^{k}$, can be represented by $\tilde{\bm{U}}_{-i}^{k}=\{\tilde{\bm{U}}_{j}^{k}\}_{j\in\mathcal{N}_{i}}$.
@@
The proposed trajectory prediction framework involves a trajectory predictor illustrated in Figure 2 and a KKT-informed training approach illustrated in Figure 3. The trajectory predictor combines the observations made by the ego UAV and the compressed messages shared by the target UAV, if any. The KKT-informed training approach facilitates training by devising a KKT-based differential layer to encode the structural information of the embedded QP into the trained neural network. The details of these components are detailed below in Sections III-AIII-B. Note that the developed learning-based model is implemented in each UAV with identical parameters.
III-A Trajectory predictor
We consider the scenario where UAV $i$ predicts the trajectory of UAV $j$ using its own observations and the messages shared by UAV $j$, if any. To this end, our trajectory prediction algorithm uses a Bayesian framework to combine the output of two modules: (i) an EvolveGCN that processes the observations of the ego UAV and yields predicted trajectories (denoted by $\tilde{\bm{U}}_{j,\text{EG}}^{k}$) and (ii) a VAE, whereby an encoder implemented on UAV $j$ compresses the planned trajectory of UAV $j$, and a decoder implemented on UAV $i$ reconstructs the compressed message to obtain the reconstructed trajectories (denoted by $\tilde{\bm{U}}_{j,\text{VAE}}^{k}$). Next, we provide the details of these two modules and the Bayesian framework.
@@ III-A Trajectory predictor
(1) EvolveGCN-based trajectory prediction module. This module enables UAV $i$ to predict the trajectories of its neighbor $j$, $\tilde{\bm{U}}_{j,\rm{EG}}^{k}$, at each time step $k$, leveraging (i) the previously predicted trajectories of UAV $\tilde{\bm{U}}_{j}^{k-1}$ at time step $k-1$, (ii) the location information within a horizon of $H$ past time steps $\bm{p}^{k,\text{hist}}=\{\bm{p}_{j}^{k-h}\}_{j\in\mathcal{N}_{i},h\in\{0,1,\cdots,H\}}$ of surrounding UAVs, measured by the onboard sensors of UAV $i$, and (iii) the geometric information of obstacles.
@@
As shown in Figure 2, the EvolveGCN (EG) architecture involves four parts: (i) a Long Short-Term Memory (LSTM)-based query module to characterize the correlations between trajectories predicted at consecutive time steps, (ii) an EvolveGCN-based interaction module to extract the spatiotemporal correlations of UAVs on the dynamic communication graph, (iii) an obstacle encoder to extract the geometric features of obstacles, and (iv) an LSTM-based prediction module to generate estimated $\tilde{\bm{U}}_{j,\text{EG}}^{k}$ and the vector of element-wise standard deviations $\bm{\Sigma}_{j,\text{EG}}^{k}$. Parts (i), (iii), and (iv) build on the trajectory planning neural network developed in [21] and extend it by allowing a probabilistic output representation for integration with VAE reconstruction outcomes. To this end, we add Part (ii) with an EvolveGCN architecture, which represents the swarm alignment by a GCN and evolves its parameters along the temporal dimension using LSTM without resorting to node embeddings. The details of the four parts are described as follows.
@@
(i) LSTM-based query module. An LSTM layer with a Fully-Connected (FC) layer is leveraged to exploit the temporal correlations between the trajectories of UAV $j$ at consecutive time steps. The activation function $\pi(\cdot)$ is chosen as ReLU for its gradient stability and computational efficiency. This encoder can be mathematically represented as:
@@
$\bm{y}_{j}^{k}=\text{Query}_{\text{LSTM}}\left(\tilde{\bm{U}}_{j}^{k-1}\right),\leavevmode\nobreak\ j\in\mathcal{N}_{i},$ (3)
@@
(ii) EG-based interaction module. As mentioned above, the EG layer is adopted to extract the spatiotemporal correlations of UAVs in time-varying communication graphs. The input to the EG layer includes the position information of $H$ past time steps $\bm{p}^{k,\rm{hist}}$ within the communication range of UAV $j$.
Each EG layer involves a GCN layer and an LSTM layer, whereby the GCN layer captures the topology of the communication graph of UAV $j$, represented as adjacency matrix $\bm{G}_{j}^{k}\in\mathbb{R}^{n\times n}$, where the elements involving unseen UAVs are set to zero. The LSTM layer is employed to model the evolution of the GCN weight matrix as a dynamical system. Specifically, let $\bm{W}^{k,(l)}$ denote the dynamic GCN weight matrix of the $l$-th layer at time step $k$, and $\bm{H}^{k,(l)}$ represents the node embeddings of the $l$-th layer at time step $k$. The EG layer can be written as:
$\bm{H}^{k+1,(l+1)}$ $=\text{GCONV}(\bm{G}_{j}^{k},\bm{H}^{k,(l)},\bm{W}^{k,(l)})$ (4a)
$\bm{W}^{k+1,(l)}$ $=\text{LSTM}_{\text{EG}}\left(\bm{W}^{k,(l)}\right).$ (4b)
where the initial embedding matrix $\bm{H}^{k,(0)}$ includes the node features as the real-time locations of surrounding UAVs and the final embedding matrix $\bm{H}^{k,(L)}$ is the output of the EG layer. In Eq.(4b), the evolution of the weight matrix $\bm{W}^{k,(l)}$ is solely based on the weight matrix at the previous time step, which is referred to as EGCU-O. An alternative way of formulating this LSTM layer is to incorporate node embeddings $\bm{H}^{k,(l)}$, i.e., $\bm{W}^{k+1,(l)}=\text{LSTM}\left(\bm{H}^{k+1,(l)},\bm{W}^{k,(l)}\right)$, referred to as EGCU-H [34]. In this paper, we use EGCU-O for its simplicity.
The output of the EG layer is further processed by a ReLu-activated FC layer to reshape the feature dimensions. The output of the interaction module, $\bm{g}^{k}$, serves as an input to the LSTM-based decoder, i.e., Part (iv).
(iii) FC-based obstacle module. Since we only consider static obstacles, we design the obstacle encoder as a fully connected layer with an input of the parameters characterizing the obstacles. This encoder can be mathematically represented as:
$\bm{o}^{k}=\textit{ReLu}\left(\text{FC}_{\text{obs}}\left(\bm{C}_{m}\right)\right),\leavevmode\nobreak\ {m\in M}$ (5)
@@
(iv) LSTM-based decoder module. The decoder synthesizes the outputs of the LSTM-based query module, the EG-based interaction module, and the FC-based obstacle module with an LSTM layer followed by two parallel fully connected layers to generate predicted trajectories of UAV $j$, $\tilde{\bm{U}}_{j}^{k}$ and the vector of element-wise standard deviations $\bm{\sigma}_{j,\text{EG}}^{k}$, which can be written as:
$\tilde{\bm{U}}_{j,\text{EG}}^{k},\bm{\sigma}_{j,\text{EG}}^{k}=\text{Decoder}_{\text{LSTM}}\left(\bm{y}_{j}^{k},\bm{o}^{k},\bm{g}^{k}\right)$ (6)
@@
where the distribution of the ground-truth trajectory is given as a Gaussian distribution with mean $\tilde{\bm{U}}_{j,\rm{EG}}^{k}$ and covariance matrix $\Sigma^{k}_{j,\text{EG}}=\left(\text{diag}(\bm{\sigma}_{j,\text{EG}}^{k})\right)^{2}$. We choose the covariance matrix as a diagonal matrix, a typical treatment in existing works to simplify the underlying neural network.
(2) VAE-based trajectory compression and reconstruction. In communication-limited scenarios, a VAE is used for trajectory compression and reconstruction. It consists of an encoder $\phi_{\rm{e}}$ and a decoder $\phi_{\rm{d}}$. The encoder is leveraged by UAV $j\in\mathcal{N}_{i}$ to compress its planned trajectory $\bm{U}_{j}^{k}\in\mathbb{R}^{3P}$ into a low-dimensional message $\bm{z}_{j}^{k}\in\mathbb{R}^{L}$ ($L<3P$). Specifically, the encoder is implemented as a neural network that outputs a vector of means $\phi_{\rm{e}}(\bm{U}_{j}^{k})\in\mathbb{R}^{L}$ and standard deviations $\bm{\sigma}_{j,\text{VAE}}^{k}\in\mathbb{R}^{L}$ of independent Gaussian distributions, from which the message $\bm{z}_{j}^{k}$ is sampled and transmitted to UAV $i$. After UAV $i$ receives the message $\bm{z}_{j}^{k}$, it uses the decoder to reconstruct the trajectory $\tilde{\bm{U}}_{j,\text{VAE}}^{k}=\phi_{\rm{d}}(\bm{z}_{j}^{k})$. In this paper, both the encoder network $\phi_{\rm{e}}$ and decoder network $\phi_{\rm{d}}$ are implemented as fully connected layers. In addition, the trajectory reconstructed by the decoder naturally makes VAE an estimator of the true planned trajectory.
@@
(3) Bayesian integration. We combine the output of the EG and VAE-based modules using a simple Bayesian approach. Specifically, we treat the outcome of the EG-based module as the prior because such prediction will always be available despite communication failure. Specifically, the prior distribution is characterized by a Gaussian distribution written as:
$\bm{U}_{j}^{k}|\tilde{\bm{U}}_{j,\text{EG}}^{k}\sim N\left(\tilde{\bm{U}}_{j,\text{EG}}^{k},\Sigma^{k}_{j,\text{EG}}\right)$ (7)
@@
where $\bm{U}_{j}^{k}$ represents the ground-truth trajectories of UAV $j$ starting from time step $k$.
@@
If UAV $i$ receives a message $\bm{z}_{j}$ from UAV $j$ and reconstructs it, the reconstructed outcome, $\tilde{\bm{U}}_{j,\text{VAE}}^{k}$, will serve as an observation, i.e.,
@@
$\tilde{\bm{U}}_{j,\text{VAE}}^{k}|\bm{U}_{j}^{k}\sim N\left(\bm{U}_{j}^{k},\Sigma^{k}_{\text{VAE}}\right),\leavevmode\nobreak\ j\in\mathcal{N}_{i}$ (8)
@@
where the covariance matrix $\Sigma^{k}_{\text{VAE}}$ can be obtained empirically from a well-trained VAE.
@@
Combining Eq.(7) and (8), the posterior distribution can be derived using the Bayesian formula as in Eq.(9). These outputs are combined using a Bayesian framework that calculates the following posterior
$p\left(\bm{U}_{j}^{k}\mid\tilde{\bm{U}}_{j,\text{VAE}}^{k},\tilde{\bm{U}}_{j,\text{EG}}^{k}\right)\propto p\left(\bm{U}_{j}^{k}\mid\tilde{\bm{U}}_{j,\text{EG}}^{k}\right)p\left(\tilde{\bm{U}}_{j,\text{VAE}}^{k}|\bm{U}_{j}^{k}\right).$ (9)
@@
where notice that $\tilde{\bm{U}}_{j,\text{VAE}}^{k}$ and $\tilde{\bm{U}}_{j,\text{EG}}^{k}$ are conditional independent. Then, the final prediction can be calculated via maximum a posteriori (MAP) as $\tilde{\bm{U}}_{j}^{k}=\arg\max_{\bm{U}_{j}^{k}}p(\bm{U}_{j}^{k}|\tilde{\bm{U}}_{j,\text{VAE}}^{k},\tilde{\bm{U}}_{j,\text{EG}}^{k})$.
III-B KKT-informed training of the trajectory prediction neural network
Sections III-A defines a learning-based framework for trajectory prediction with a neural network model. One classical method to train the neural network is to minimize the $l_{2}$-loss between the true planned trajectories $\bm{U}_{j}^{k}$ and the predicted trajectories $\tilde{\bm{U}}_{j}^{k}$ for each UAV $j$ at each time step $k$, written as
$\mathcal{L}_{2}(\tilde{\bm{U}}_{j}^{k},\bm{U}_{j}^{k})=||\tilde{\bm{U}}_{j}^{k}-\bm{U}_{j}^{k}||_{2}^{2}$ (10)